OpenStudy (anonymous):

Solve the equation on the interval [0, 2π). cos x + 2 cos x sin x = 0 A. 0 ; π/2 ; 7π/6 ; 3π/2 B. π/2 ; 7π/6 ; 3π/2 ; 11π/6 C. 7π/6 ; 11π/6 ; 2n D. 7π/6 ; 11π/6

OpenStudy (anonymous):

Do you need step by step help or just guidance?

OpenStudy (anonymous):

\[\cos(x)(1+2\sin(x))=0\] \[\cos(x)=0\implies x=\frac{\pi}{2},x=\frac{3\pi}{2}\]or \[1+\sin(x)=0\implies \sin(x)=\frac{1}{2}\]etc

OpenStudy (anonymous):

would it be B

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thank you very much. i just get stuck in the first steps.

OpenStudy (anonymous):

Ok, np.